OpenStudy (anonymous):

need to find m so that y=mx+5 has a common intersection point with y=7x+43 and y=8x+53

OpenStudy (anonymous):

Do you mean same common intersection point

OpenStudy (anonymous):

If no all values of m except 7,8

OpenStudy (anonymous):

i need to find a value for the slope for thee first equation so it has a common intersection point with the other 2

OpenStudy (anonymous):

m=16/5

OpenStudy (anonymous):

unfortunately that didn't solve it

OpenStudy (mr.math):

Do you know how to find the intersection point of the two given lines?

OpenStudy (mr.math):

I mean y=7x+43 and y=8x+53.

OpenStudy (mr.math):

The two lines intersect \(7x+43=8x+53 \implies x=-10\). Substitute \(x=-10\) in either equation you get \(y=-70+43=-27\). So the intersection point is \(-10,-27\). Plug this point in the first equation and solve for m.

OpenStudy (anonymous):

Mr Math says the same

OpenStudy (anonymous):

I did it using determinats

OpenStudy (mr.math):

Yes, NotSObright is so bright and he's right.

OpenStudy (mr.math):

You basically have to solve: \[-27=-10m+5 \implies m=\frac{32}{10}=\frac{16}{5}.\]

OpenStudy (anonymous):

my mistake. Thank you to both of you. The online program I use is picky... did not have to simplify to 16/5. successful using 32/10